OpenStudy (anonymous):

John has four times as many marbles as Neil. If John gave 18 to Neil, they would have the same number. How many marbles has each??

OpenStudy (anonymous):

Number of marbles of Neil = x Number of marbles of john=4x 4x-18=x+18 3x=36 x=12 Neil=12 John=48

OpenStudy (anonymous):

brilliant, I see where I was going wrong. thank you so much.

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

why did you let x=neil??

OpenStudy (anonymous):

Because it is given in the question that John has "four times" as many marbles as "neil"

OpenStudy (anonymous):

So let neil be x & john 4 times x = 4x

OpenStudy (anonymous):

nice.

OpenStudy (anonymous):

Let John be J and Neil be N, 1) "John has four times as many marbles as Neil" i.e J=4N 2)"If John gave 18 to Neil" i.e J=18+N now solve by substituting the value for J in the 1st statement , so the equation will be J=18+N 4N=18+N (J=4N, from the ist statement) 4N-N=18 3N=18 N=6 Therefore J=4*6 i.e 24 and since John gave 18 to Neil, 18+6=24.

OpenStudy (anonymous):

So if John has 24 and Neil has 6 ,When John gives 18 to Neil ,John's marbles reduces to 6 and Neil's increases to 24 hence Both of them does not have the same number . So i assume your answer is wrong.

OpenStudy (anonymous):

Thank you trying nidi.

OpenStudy (anonymous):

not a problem :-) i'm looking at the question stem again.... and reading it in parts..

OpenStudy (anonymous):

Nidi You made a mistake in this step J=18+N When John give 18 Marbles it will be J-18=18+N Then you can continue with your method :) Good one

OpenStudy (anonymous):

oooopsss!!!! yes that's where i went wrong !!! apologies!! :P

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

john's marbles=4(Neil's marbles),let x=john's marbles;y=Neil's marbles x=4y; x-18=y+18; 4y-18=y+18; 3y=36; y=12-Neils'smarbles x=4y; x=4(12) x=48-john's marbles

OpenStudy (anonymous):

nice